OpenStudy (anonymous):

Anyone please help and I will fan and give medals!!!! I need someone to check if I'm right!Graph the system of inequalities. Then use your graph to identify the point that represents a solution to the system.

OpenStudy (anonymous):

@wendylisette

OpenStudy (anonymous):

x+y is greater than or equal to 3 x-3y<2 A.(6,1) B. (8-1) C. (6,2) D.(6,-2)

OpenStudy (anonymous):

@texaschic101 please help!!!!

OpenStudy (anonymous):

I think its C.

OpenStudy (texaschic101):

(6,1) x = 6 and y = 1 x + y >= 3 x - 3y < 2 6 + 1 >= 3 6 - 3(1) < 2 7 >= 3 (correct) 6 - 3 < 2 3 < 2 (false) not the first one (-8,1) x + y >= 3 -8 + 1 >= 3 -7 >= 3 (false) No need to check the other equation because the points have to satisfy both equations (6,2) x + y >= 3 x - 3y < 2 6 + 2 >= 3 6 - 3(2) < 2 8 >= 3 (correct) 6 - 6 < 2 0 < 2 (correct) (6,2) -- answer C, satisfies both equations

OpenStudy (anonymous):

omg thank you so much!!!!!!

OpenStudy (texaschic101):

no problem :)

OpenStudy (anonymous):

can you another one?

OpenStudy (texaschic101):

sure :)

OpenStudy (anonymous):

3x-4y>0 x-5y>0

OpenStudy (anonymous):

A. (0,0) B. (6,0) C. (6,2) D. (-4,-2) I think its B.

OpenStudy (texaschic101):

(6,0) 3x - 4y > 0 x - 5y > 0 3(6) - 4(0) > 0 6 - 5(0) > 0 18 > 0 (true) 6 > 0 (correct) answer is (6,0)

OpenStudy (anonymous):

YAY thank you so much I will probably bug you in the near future for help since your so great at explaining thank you soooo much!!!!! :)

OpenStudy (texaschic101):

glad to help :)